# Content ARCs: Decentralized Content Rights in the Age of Generative AI

### v1

@@ Abstract
1 Purpose
Generative AI (GenAI) is transforming creative workflows through tools for efficiently generating and manipulating text, images, music, and video. However, the vast datasets used to train commercial GenAI models often include copyrighted material, causing growing concerns among creative rightsholders about how their work is used and whether they will receive proper recognition and compensation. These concerns have recently surfaced in proposals—and counter-proposals—around changes to copyright legislation in several jurisdictions. The debate has become highly contested; for example, the recent UK copyright consultation [27] received over 11,000 submissions following a public awareness campaign by creative practitioners. Arguments often centre on the contentious issue of opt-in versus opt-out mechanisms; i.e. whether permission should always be sought to train commercial AI models on copyrighted data (i.e. opt-in) or there should be a default legal right to train unless counter-indicated by a rightsholder (i.e. opt-out). Rightsholders often argue that an opt-in model is essential to protect creators’ control over their work and ensure fair compensation [16]. While, AI developers often contend that an opt-out approach is necessary given the sheer scale of data (often many billions of assets) over which contractual consent and payment would need to be orchestrated [17]. Political factors also weigh on the debate, such as creating a competitive AI development environment versus jurisdictions with more permissive legislation or less stringent enforcement [12]. These positions are often deeply polarized and appear difficult to reconcile, highlighting the need for balanced solutions that uphold both creative rights and technological progress.
@@ 3.3 Compensation
Attribution-Driven Payouts. The training datasets responsible for creating AI models can be described via provenance metadata attached to models, raising the opportunity to pay royalties to contributors for use of the model e.g. when synthetic assets are generated [11]. For Generative AI the scale of such datasets runs into the billions, presenting a challenge to fractional compensation. Data attribution can help identify the subset of training data most influential in a given synthetic asset, to target payout. Several techniques have measured attribution as visual correlation between output and training data [6, 30]. Yet correlation does not equal causation. Invisible watermarking of training data has been explored for causative attribution in ProMark[5] to measure which mixture of watermarks exists in the output, but does not yet scale to billions of data items. Analytic approaches like Shapley values [29] are also causative but have yet to be demonstrated at a scale needed for generative AI.
@@ 4 Discussion
5 Limitations

### v2

@@ Abstract
1 Introduction
Generative AI (GenAI) is transforming creative workflows through tools for efficiently generating and manipulating text, images, music, and video. However, the vast datasets used to train commercial GenAI models often include copyrighted material [1], causing growing concerns among creative rightsholders about how their work is used and whether they will receive proper recognition and compensation. These concerns have recently surfaced in proposals—and counter-proposals—around changes to copyright legislation in several jurisdictions. The debate has become highly contested; for example, the recent UK copyright consultation [2] received over 11,000 submissions following a public awareness campaign by creative practitioners. Arguments often centre on the contentious issue of opt-in versus opt-out mechanisms; i.e. whether permission should always be sought to train commercial AI models on copyrighted data (i.e. opt-in) or there should be a default legal right to train unless counter-indicated by a rightsholder (i.e. opt-out). Rightsholders often argue that an opt-in model is essential to protect creators’ control over their work and ensure fair compensation [3]. While, AI developers often contend that an opt-out approach is necessary given the sheer scale of data (often many billions of assets) over which contractual consent and payment would need to be orchestrated [4]. Political factors also weigh on the debate, such as creating a competitive AI development environment versus jurisdictions with more permissive legislation or less stringent enforcement [5]. These positions are often deeply polarized and appear difficult to reconcile, highlighting the need for balanced solutions that uphold both creative rights and technological progress.
@@ 3.2 Rights
While provenance has seen increasing standardisation through initiatives like C2PA, no equivalent effort yet exists for machine-readable rights and compensation. We highlight the need for coordinated standardisation across these phases to enable end-to-end interoperability and scalable, decentralised licensing for content.
@@ 3.3 Compensation
Attribution-Driven Payouts. The training datasets responsible for creating AI models can be described via provenance metadata attached to models, raising the opportunity to pay royalties to contributors for use of the model e.g. when synthetic assets are generated [18]. However, for Generative AI the scale of such datasets runs into the billions, presenting a challenge to fractional compensation. Data attribution can help identify the subset of training data most influential in a given synthetic asset, to target payout. Several techniques have measured attribution as visual correlation between output and training data [28, 29]. Yet correlation does not equal causation. Causative methods have been proposed, such as invisible watermarking in ProMark[30] to measure which mixture of watermarks exists in the output, as well as Shapley value-based analytics [31], however neither scales to billions of data items. Scaling causative attribution techniques to the billion-scale datasets used for training GenAI models remains an open research challenge in the development of royalty mechanisms for GenAI.
@@ 4 Discussion
5 Conclusion
Once implemented, Content ARCs would enable end-to-end machine-readable permissions and automatic compensation throughout the AI development. A future case study could involve a content creator utilizing ARC-compliant tools to manage rights and receive compensation for their work used in AI training. This demonstrates how our framework directly addresses current policy needs, responding to the needs outlined in the UK’s recent consultation on AI and copyright, which addresses transparent licensing and fair remuneration for creators in AI development [2].